OpenStudy (anonymous):

What equations are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1) ?

OpenStudy (anonymous):

Find the slope first.. Slope = 2 Now, the parallel line equation is: y - y1 = m(x - x1) y - 1 = 2(x - 4) y = 2x - 7 Now for perpendicular lines, Slope = negative reciprocal of given slope = -1/2 Now use the same equation to find its equation: y - y1 = m(x - x1) Here put m = -1/2 And find the equation..

OpenStudy (anonymous):

Would the equation be : parallel: y = 2x + 2 perpendicular: y = x +

OpenStudy (anonymous):

parallel y = 2x + 7 perpendicular y = -x/2 + b substitute x = 4, y = 1 to find b